seperate 43 people into two groups so that the first group is 5 less than three times the second.
I don't really know how to describe the criteria of this problem, but it surely has something to do with algebra.

Answer by Paul(988)

Let x be first group and let 3x-5 be the second
x+(3x-5)=43
4x-5=43
4x-5+5=43+5
4x=48
4x/4 = 48/4
x=12
3x12-5 = 31
Therefore there is one group of 12 and the other of 31